OpenStudy (maddy1251):

Need help with a Conic Sections equation? Identify the Center and Radius for the following equation:

OpenStudy (maddy1251):

\[x^{2}+y^{2}=7\]

OpenStudy (maddy1251):

I believe the radius to be about 2.6.

imqwerty (imqwerty):

center =(0,0) radius =sqrt{7}

imqwerty (imqwerty):

yes sqrt{7}=2.6

OpenStudy (maddy1251):

Thank you @imqwerty ! I didn't know what to do with the (x,y) values. Much thanks :-)

imqwerty (imqwerty):

(: np did u get the center?

OpenStudy (maddy1251):

@imqwerty yeah :) I did! I kinda thought it would be either (1,1) or (0,0), just wasn't 100% which it would be

imqwerty (imqwerty):

ok when ever u have this kinda thing compare it with equation of circle equation of circle-\[(x-h)^2 +(y-k)^2=r^2\] where (h.k)=center coordinates r=radius of cricle

OpenStudy (johnweldon1993):

Just remember the equation of a circle \[\large (x - h)^2 + (y - k)^2 = r^2\] (h,k) is your center

OpenStudy (johnweldon1993):

Oh beat me to it haha

imqwerty (imqwerty):

lol

OpenStudy (maddy1251):

@imqwerty and @johnweldon1993 thanks :) I have that formula written down, we are just now starting this stuff.

imqwerty (imqwerty):

np (B

OpenStudy (johnweldon1993):

Of course, and don't worry...they don't get too bad :D

OpenStudy (maddy1251):

@johnweldon1993 Hopefully not.. xD